OpenStudy (anonymous):

Let x = 2 sin θ, − π/2 < θ < π/2. Simplify the expression. x/square root of 4-x^2

OpenStudy (anonymous):

ok i get it it means replace \(x\) by \(2\sin(\theta)\) right?

OpenStudy (anonymous):

\[\frac{x}{\sqrt{4-x^2}}\] first make the direct replacement and get \[\frac{2\sin(\theta)}{\sqrt{4-(2\sin(\theta)^2}}\]

OpenStudy (anonymous):

then square and get \[\frac{2\sin(\theta)}{\sqrt{4-4\sin^2(\theta)}}\]

OpenStudy (anonymous):

factor out the 4 and get \[\frac{2\sin(\theta)}{\sqrt{4(1-\sin^2(\theta))}}\]

OpenStudy (anonymous):

then since \(1-\sin^2(\theta)=\cos^2(\theta)\) you have \[\frac{2\sin(\theta)}{\sqrt{4\cos^2(\theta)}}=\frac{2\sin(\theta)}{2\cos(\theta)}=\tan(\theta)\]

OpenStudy (anonymous):

Thank you :)